# The role of peculiar velocity uncertainties in standard siren cosmology

### v1

@@ Abstract
Local distance indicators such as standard sirens, in combination with spectroscopic redshift measurements of their host galaxies, allow us to estimate the present-day expansion rate of the Universe parameterised by Hubble’s constant, $H_{0}$. However, these observed redshifts are systematically modified by the effect of galaxy peculiar velocities. Although these velocities may be estimated from the local density field by the process of velocity-field reconstruction, the intrinsic errors and covariance in these estimates contribute to the error in the $H_{0}$ determination. In this paper we demonstrate how the impact of peculiar velocities can be propagated into $H_{0}$ measurements from local distance indicators with observed redshifts, incorporating the full covariance of the velocity field induced by bulk flows. We apply our methods to cosmological simulations, testing the importance of this effect in the context of future analyses of gravitational wave sources with electromagnetic counterparts used as bright sirens. We conclude that $H_{0}$ errors may be increased by up to a factor of 2 in comparison with neglecting peculiar velocity covariance, with the highest impacts expected for sources at nearby distances or with small distance errors. Our analytical methods may also be applied to other local distance indicators, such as Type Ia supernovae.
@@ 1 Introduction
The present-day expansion rate of the Universe, known as Hubble’s constant $H_{0}$, is a key parameter of our cosmological model, linking distance and redshift in the local Universe. Its value may be measured by combining the recession velocity of nearby objects with redshift-independent distance indicators, such as standard candles. However, current determinations of $H_{0}$ from local “distance ladder” techniques (e.g., Riess et al., 2016; Freedman et al., 2019; Pesce et al., 2020; Blakeslee et al., 2021; Riess et al., 2022; Said et al., 2025; Freedman et al., 2025) lie in moderate-to-significant disagreement with predictions of the value of $H_{0}$ based on the standard cosmological model, calibrated by observations of the Cosmic Microwave Background radiation (Planck Collaboration et al., 2020; Balkenhol et al., 2023; Madhavacheril et al., 2024) and Baryon Acoustic Oscillations (Beutler et al., 2011; Alam et al., 2021; DESI Collaboration et al., 2025). This disagreement is known as the “Hubble tension”, and has motivated a wide range of potential solutions (see reviews by, e.g., Mörtsell & Dhawan, 2018; Di Valentino et al., 2021; Freedman, 2021; Vagnozzi, 2023), many of which challenge the assumptions of the current cosmological model.
@@
Gravitational-wave astronomy has enabled a new type of distance indicator for mapping out local expansion: standard sirens, which result from mergers of compact binary systems of black holes or neutron stars. By fitting binary coalescence models to the observed gravitational-wave signals, the physical parameters of each merger may be determined, and hence the amplitude of the signal can be used to infer the distance to the event. If the gravitational-wave source has a known electromagnetic counterpart (“bright siren”), whose redshift may also be measured, the Hubble constant may be directly determined. For example, the first bright siren event, GW170817, produced a constraint $H_{0}=70^{+12}_{-8}$ km s-1 Mpc-1 (Abbott et al., 2017), which has been refined by several later studies (e.g., Hotokezaka et al., 2019; Howlett & Davis, 2020; Nicolaou et al., 2020; Palmese et al., 2024; Gourdji et al. in prep., 2025). In the absence of electromagnetic counterparts, gravitational-wave sources may be associated with the distribution of galaxies in a statistical sense as “dark sirens” (Schutz, 1986; Soares-Santos et al., 2019; Palmese et al., 2020; Gray et al., 2023; Gair et al., 2023; Alfradique et al., 2024; Cross-Parkin et al., 2025). Whilst the accuracy of standard siren measurements is not currently competitive with other late-Universe determinations of $H_{0}$, the number of available systems, and the precision of the resulting distance measurements, will improve over coming years, with forecasts presented by, e.g., Chen et al. (2018); Mortlock et al. (2019); Feeney et al. (2019).
@@
A key challenge when determining the Hubble constant from local distance indicators is that the observed redshifts of objects reflect not just the overall cosmological expansion, but also the local motions or “peculiar velocities” of galaxies that result from gravitationally-driven flows towards overdense regions (for a recent review, see Turner, 2024). Moreover, the peculiar velocity has the highest fractional contribution to the redshift for the nearby sources which provide the most accurate distance measurements. As such, peculiar velocities contribute a potentially important source of statistical and systematic error to $H_{0}$ determinations. The connection between peculiar velocities and standard siren inferences has been considered by various studies (Wang et al., 2018; He, 2019; Howlett & Davis, 2020; Nicolaou et al., 2020; Mukherjee et al., 2021; Nimonkar & Mukherjee, 2024). As a simple measure, a typical peculiar velocity error range might be marginalized. Alternatively, the velocity field at the locations of the distance indicators may be “reconstructed” from the overall galaxy distribution using linear-theory models or a variety of algorithms (e.g., Dekel et al., 1993; Nusser & Davis, 1994; Zaroubi et al., 1999; Kitaura et al., 2012; Carrick et al., 2015; Lavaux, 2016; Jasche & Lavaux, 2019; Lilow & Nusser, 2021; Qin et al., 2023), and used to correct the observed redshifts, albeit with some margin of error.
@@
An issue generally neglected in previous studies is that both the underlying and reconstructed peculiar velocities of objects are correlated between different locations (Gorski, 1988; Blake & Turner, 2024a), because the peculiar velocity field is organised in large-scale “bulk flows” on $\sim 100$ Mpc scales. In this paper we investigate the impact of these velocity correlations on determinations of $H_{0}$ from local distance indicators, in the context of simulated bright siren analyses. We assess the extent to which these correlated velocity errors affect the error in the $H_{0}$ determination as a function of the number of bright sirens, their distance range, the error in the inferred distances and the inclusion of velocity reconstruction. Our study hence extends the analysis of Turner & Blake (2023) to include velocity covariance, and should provide a useful starting point for the treatment of correlated peculiar velocity errors in future analyses of standard sirens and other distance indicators.
@@ 2 Velocity-field statistics
In standard-siren cosmology, we wish to infer the value of Hubble’s constant $H_{0}$ from determinations of the distance $D$ and redshift $z$ of sources. For the purposes of our analysis and clarity of presentation, we assume a low-redshift approximation in which $D=cz_{\rm cos}/H_{0}$. Whilst we caution that this approximation is not appropriate for analysis of real datasets (as discussed by Davis et al., 2019), this model is applied self-consistently in our simulated surveys described below, and therefore does not limit this work. The observed redshift $z$ is a combination of the cosmological redshift $z_{\rm cos}$ and the radial peculiar velocity $\mathrm{u}$,
@@
The peculiar velocity field may be estimated from the observed galaxy density field by a process of “velocity-field reconstruction”, and then applied in Eq. 1 to “correct” the observed redshifts and improve the determination of $H_{0}$. However, the reconstructed velocity field is subject to error due to sample variance, noise (from the limited number density of galaxies) and non-linear effects. Turner & Blake (2023) argued that a useful statistical model for the joint probability of the true underlying velocity $\mathrm{u}(\mathbf{x})$ and the reconstructed velocity $\mathrm{v}(\mathbf{x})$ at a location $\mathbf{x}$ is a correlated Gaussian likelihood distribution,
@@
in terms of variances $\sigma_{\mathrm{v}}^{2}=\langle\mathrm{v}^{2}(\mathbf{x})\rangle$, $\sigma_{\mathrm{u}}^{2}=\langle\mathrm{u}^{2}(\mathbf{x})\rangle$, and the cross-correlation coefficient $r=\sigma_{\mathrm{u}\mathrm{v}}^{2}/(\sigma_{\mathrm{u}}\,\sigma_{\mathrm{v}})$, where $\sigma_{\mathrm{u}\mathrm{v}}^{2}=\langle\mathrm{u}(\mathbf{x})\,\mathrm{v}(\mathbf{x})\rangle$. Expressions for $\sigma_{\mathrm{u}}^{2}$, $\sigma_{\mathrm{v}}^{2}$ and $\sigma_{\mathrm{u}\mathrm{v}}^{2}$ are given by Blake & Turner (2024a) for different types of velocity-field reconstruction. In this statistical model, given the value of the reconstructed velocity $\mathrm{v}(\mathbf{x})$ at a point, the posterior distribution of the underlying velocity $\mathrm{u}(\mathbf{x})$ is a Gaussian with mean $(r\,\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}})\,\mathrm{v}$ and variance $(1-r^{2})\,\sigma_{u}^{2}$.
The values of the underlying and reconstructed velocities are correlated between different locations, because the velocity field is arranged in “bulk flows” on $\sim 100$ Mpc scales. The form of these covariances depends on the type of velocity-field reconstruction applied (Blake & Turner, 2024a), and will be discussed further in Sec. 4. These correlations have potentially important implications for joint analyses of multiple standard sirens.
@@ 3 Determining the $H_{0}$ posterior
$P(\mathrm{u})\propto\exp{\left[-\frac{\left(c\,z-H_{0}\,D-(r\,\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}})\,\mathrm{v}\right)^{2}}{(1-r^{2})\,\sigma_{u}^{2}}\right]}.$ (5)
@@
and the covariance of the velocity differences $\mathbf{C}_{\mathrm{w}}$ is evaluated using Eq. 32 in Blake & Turner (2024a), and will be described in more detail in Sec. 4 for the velocity reconstruction method we adopt in our current study.
If we represent the luminosity distance probability distributions $P(D_{\alpha})$ as Gaussian distributions with mean $\mu_{D,\alpha}$ and standard deviation $\sigma_{D,\alpha}$, Eq. 6 can be conveniently evaluated by analytical marginalisation techniques (Bridle et al., 2002; Taylor & Kitching, 2010). Following the notation of Taylor & Kitching (2010), we expand the log-likelihood $\mathcal{L}=-2\ln{P(\mathrm{u}_{\alpha})}$ of the inferred peculiar velocity field values as a Taylor series as a function of $\delta D_{\alpha}=D_{\alpha}-\mu_{D,\alpha}$,
@@
The marginalised log-likelihood $\mathcal{L}_{\rm marg}=-2\ln{L_{\rm marg}}$, after integrating over the Gaussian priors in $D_{\alpha}$, is then given by Eq. 15 in Taylor & Kitching (2010),
@@ 4 Simulation
We tested our formalism for $H_{0}$ recovery by constructing a model bright siren analysis of a simulated dataset. For this purpose we used the $z=0$ snapshot of the Gigaparsec WiggleZ N-body (GiggleZ) simulation (Poole et al., 2015), which is a dark matter-only simulation run in a $1\,h^{-1}$ Gpc periodic box with a halo mass resolution of $3.0\times 10^{11}h^{-1}M_{\odot}$. This simulation was created using the fiducial WMAP-5 cosmological parameters: $\Omega_{m}=0.273$, $\Omega_{\Lambda}=0.727$, $\Omega_{b}=0.0456$, $h=0.705$, $\sigma_{8}=0.812$ and $n_{s}=0.960$.
@@
  • We added a Gaussian error to the distances equal to a constant fraction of the true distances, $\sigma_{D,\alpha}=f_{D}\,D_{{\rm true},\alpha}$, forming the noisy distances, $D_{\alpha}$.
@@
We repeated this analysis varying the number of bright sirens $N$, their maximum distance $D_{\rm max}$, and the fractional distance error $f_{D}$.
For convenience of this analysis we assume a “Fourier-based” linear-theory velocity reconstruction approach, which uses a density-field defined within a Fourier cuboid. Blake & Turner (2024a) contrast the Fourier-based approach with the more computationally-intensive spherical Fourier-Bessel reconstruction within a curved-sky window function, demonstrating in each case how the sample variance and noise propagates into the statistics of the reconstructed velocity field. However, in our study we focus instead on the propagation of the velocity field statistics into the posterior probability distribution of $H_{0}$. In this case, we simply assume a uniform selection function within the box.
@@
In this study we focus on the radial velocity of each source relative to an observer at the origin, $\mathrm{u}(\mathbf{x})=\sum_{i}\mathrm{u}_{i}(\mathbf{x})\,x_{i}/x$. The correlation between the radial velocities at two locations is given from Eq. 13 by (e.g., Gorski, 1988; Blake & Turner, 2024b),
@@
where $j_{\ell}$ are the spherical Bessel functions of order $\ell$ and $P_{m}(k)$ is the matter power spectrum, which we obtain for the GiggleZ fiducial cosmology using the CAMB and halofit prescriptions (Lewis et al., 2000; Smith et al., 2003; Takahashi et al., 2012).
@@
Following Blake & Turner (2024a), the covariance of the reconstructed velocities between two positions, $\langle\mathrm{v}(\mathbf{x})\,\mathrm{v}(\mathbf{y})\rangle$, is obtained using the same relation given in Eq. 14 whilst making the replacement,
@@
These covariances allow us to determine the covariance of velocity differences $\mathbf{C}_{\mathrm{w}}$ introduced in Eq. 7, using Eq. 32 in Blake & Turner (2024a):
@@
where the factor $R=r\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}}$ and $\epsilon_{\mathrm{u}}$ is the additional noise in the underlying velocities due to, for example, non-linear effects. We assume $\epsilon_{\mathrm{u}}=150\,{\rm km}\,{\rm s}^{-1}$ for the purposes of this study, which is a standard choice for residual velocity noise (Carrick et al., 2015; Scolnic et al., 2018). Hence, we distinguish two contributions to the velocity errors: linear-theory covariance (due to the imperfection of velocity reconstruction, represented by the first four terms of Eq. 17) and non-linear noise (represented by the final term). For our fiducial assumptions the values of the coefficients are given by $\sigma_{u}=337\,{\rm km}\,{\rm s}^{-1}$, $\sigma_{v}=220\,{\rm km}\,{\rm s}^{-1}$ and $r=0.72$.
@@
In the following section we compare the results of applying these different analysis approaches for various choices of $N$, the number of sirens, $D_{\rm max}$, the maximum observed distance, and $f_{D}$, the fractional error in those distance measurements. We considered $N=[2,5,10,25,100]$, $D_{\rm max}=[50,100,150,200]$ Mpc, and $f_{D}=[0.01,0.02,0.05,0.10]$ to define this parameter space, making for 80 total tests. We repeated each of these tests for 50 different sub-samples drawn from the N-body simulation, saving the mean and standard deviation obtained from the $H_{0}$ posterior for each realisation. We then use the mean of these sets of results for a final measurement for each configuration.

### v2

@@ Abstract
Local distance indicators such as standard sirens, in combination with spectroscopic redshift measurements of their host galaxies, allow us to estimate the present-day expansion rate of the Universe parameterised by Hubble’s constant, $H_{0}$. However, these observed redshifts are systematically modified by the effect of galaxy peculiar velocities. Although these velocities may be estimated from the local density field by the process of velocity-field reconstruction, the intrinsic errors and covariance in these estimates contribute to the error in the $H_{0}$ determination. In this paper we demonstrate how the impact of peculiar velocities can be propagated into $H_{0}$ measurements from local distance indicators with observed redshifts, incorporating the full covariance of the velocity field induced by bulk flows. We apply our methods to cosmological simulations, testing the importance of this effect in the context of future analyses of gravitational wave sources with electromagnetic counterparts used as bright sirens. We conclude that $H_{0}$ errors may be increased by a factor of 2 in comparison with neglecting peculiar velocity covariance, for GW170817-like sirens located within 50 Mpc with $\sim 5\%$ distance errors, with the highest impacts expected for sources at nearby distances or with small distance errors. Our analytical methods may also be applied to other local distance indicators, such as Type Ia supernovae.
@@ 1 Introduction
The present-day expansion rate of the Universe, known as Hubble’s constant $H_{0}$, is a key parameter of our cosmological model, linking distance and redshift in the local Universe. Its value may be measured by combining the recession velocity of nearby objects with redshift-independent distance indicators, such as standard candles. However, current determinations of $H_{0}$ from local “distance ladder” techniques (e.g., 2016ApJ...826...56R; 2019ApJ...882...34F; 2020ApJ...891L...1P; 2021ApJ...911...65B; 2022ApJ...934L...7R; 2025MNRAS.539.3627S; 2025ApJ...985..203F) lie in moderate-to-significant disagreement with predictions of the value of $H_{0}$ based on the standard cosmological model, calibrated by observations of the Cosmic Microwave Background radiation (2020A&A...641A...6P; 2023PhRvD.108b3510B; 2024ApJ...962..113M) and Baryon Acoustic Oscillations (2011MNRAS.416.3017B; 2021PhRvD.103h3533A; 2025JCAP...02..021A). This disagreement is known as the “Hubble tension”, and has motivated a wide range of potential solutions (see reviews by, e.g., 2018JCAP...09..025M; 2021CQGra..38o3001D; 2021ApJ...919...16F; 2023Univ....9..393V), many of which challenge the assumptions of the current cosmological model.
@@
Gravitational-wave astronomy has enabled a new type of distance indicator for mapping out local expansion: standard sirens, which result from mergers of compact binary systems of black holes or neutron stars. By fitting binary coalescence models to the observed gravitational-wave signals, the physical parameters of each merger may be determined, and hence the amplitude of the signal can be used to infer the distance to the event. If the gravitational-wave source has a known electromagnetic counterpart (“bright siren”), whose redshift may also be measured, the Hubble constant may be directly determined. For example, the first bright siren event, GW170817, produced a constraint $H_{0}=70^{+12}_{-8}$ km s-1 Mpc-1 (2017Natur.551...85A), which has been refined by several later studies (e.g., 2019NatAs...3..940H; 2020MNRAS.492.3803H; 2020MNRAS.495...90N; 2024PhRvD.109f3508P; KellyPaper). In the absence of electromagnetic counterparts, gravitational-wave sources may be associated with the distribution of galaxies in a statistical sense as “dark sirens” (1986Natur.323..310S; 2019ApJ...876L...7S; 2020ApJ...900L..33P; 2023JCAP...12..023G; 2023AJ....166...22G; 2024MNRAS.528.3249A; 2025arXiv250217747C). Whilst the accuracy of standard siren measurements is not currently competitive with other late-Universe determinations of $H_{0}$, the number of available systems, and the precision of the resulting distance measurements, will improve over coming years, with forecasts presented by, e.g., 2018Natur.562..545C; 2019PhRvD.100j3523M; 2019PhRvL.122f1105F.
@@
A key challenge when determining the Hubble constant from local distance indicators is that the observed redshifts of objects reflect not just the overall cosmological expansion, but also the local motions or “peculiar velocities” of galaxies that result from gravitationally-driven flows towards overdense regions (for a recent review, see 2024arXiv241119484T). Moreover, the peculiar velocity has the highest fractional contribution to the redshift for the nearby sources which provide the most accurate distance measurements. As such, peculiar velocities contribute a potentially important source of statistical and systematic error to $H_{0}$ determinations (see, e.g., 2014PhRvL.112v1301B; 2021PhRvD.104h3506F; 2023MNRAS.526..337T). The connection between peculiar velocities and standard siren inferences has been considered by various studies (2018PhRvD..98f3503W; 2019PhRvD.100b3527H; 2020MNRAS.492.3803H; 2020MNRAS.495...90N; 2021A&A...646A..65M; 2024MNRAS.527.2152N). As a simple measure, a typical peculiar velocity error range might be marginalized. Alternatively, the velocity field at the locations of the distance indicators may be “reconstructed” from the overall galaxy distribution using linear-theory models or a variety of algorithms (e.g., 1993ApJ...412....1D; 1994ApJ...421L...1N; 1999ApJ...520..413Z; 2012MNRAS.427L..35K; 2015MNRAS.450..317C; 2016MNRAS.457..172L; 2019A&A...625A..64J; 2021MNRAS.507.1557L; 2023JCAP...06..062Q; 2025arXiv250200121S), and used to correct the observed redshifts, albeit with some margin of error.
@@
An issue generally neglected in previous studies is that both the underlying and reconstructed peculiar velocities of objects are correlated between different locations (1988ApJ...332L...7G; 2024OJAp....7E..87B), because the peculiar velocity field is organised in large-scale “bulk flows” on $\sim 100$ Mpc scales. In this paper we investigate the impact of these velocity correlations on determinations of $H_{0}$ from local distance indicators, in the context of simulated bright siren analyses. We assess the extent to which these correlated velocity errors affect the error in the $H_{0}$ determination as a function of the number of bright sirens, their distance range, the error in the inferred distances and the inclusion of velocity reconstruction. Our study hence extends the analysis of 2023MNRAS.526..337T to include velocity covariance, and should provide a useful starting point for the treatment of correlated peculiar velocity errors in future analyses of standard sirens and other distance indicators.
@@ 2 Velocity-field statistics
In standard-siren cosmology, we wish to infer the value of Hubble’s constant $H_{0}$ from determinations of the distance $D$ and redshift $z$ of sources. For the purposes of our analysis and clarity of presentation, we assume a low-redshift approximation in which $D=cz_{\rm cos}/H_{0}$. Whilst we caution that this approximation is not appropriate for analysis of real datasets (as discussed by 2019MNRAS.490.2948D), this model is applied self-consistently in our simulated surveys described below, and therefore does not limit this work. The observed redshift $z$ is a combination of the cosmological redshift $z_{\rm cos}$ and the radial peculiar velocity $\mathrm{u}$,
@@
The peculiar velocity field may be estimated from the observed galaxy density field by a process of “velocity-field reconstruction”, and then applied in Eq. 1 to “correct” the observed redshifts and improve the determination of $H_{0}$. However, the reconstructed velocity field is subject to error due to sample variance, noise (from the limited number density of galaxies) and non-linear effects. 2023MNRAS.526..337T argued that a useful statistical model for the joint probability of the true underlying velocity $\mathrm{u}(\mathbf{x})$ and the reconstructed velocity $\mathrm{v}(\mathbf{x})$ at a location $\mathbf{x}$ is a correlated Gaussian likelihood distribution,
@@
in terms of variances $\sigma_{\mathrm{v}}^{2}=\langle\mathrm{v}^{2}(\mathbf{x})\rangle$, $\sigma_{\mathrm{u}}^{2}=\langle\mathrm{u}^{2}(\mathbf{x})\rangle$, and the cross-correlation coefficient $r=\sigma_{\mathrm{u}\mathrm{v}}^{2}/(\sigma_{\mathrm{u}}\,\sigma_{\mathrm{v}})$, where $\sigma_{\mathrm{u}\mathrm{v}}^{2}=\langle\mathrm{u}(\mathbf{x})\,\mathrm{v}(\mathbf{x})\rangle$. Expressions for $\sigma_{\mathrm{u}}^{2}$, $\sigma_{\mathrm{v}}^{2}$ and $\sigma_{\mathrm{u}\mathrm{v}}^{2}$ are given by 2024OJAp....7E..87B for different types of velocity-field reconstruction. In this statistical model, given the value of the reconstructed velocity $\mathrm{v}(\mathbf{x})$ at a point, the posterior distribution of the underlying velocity $\mathrm{u}(\mathbf{x})$ is a Gaussian with mean $(r\,\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}})\,\mathrm{v}$ and variance $(1-r^{2})\,\sigma_{u}^{2}$.
The values of the underlying and reconstructed velocities are correlated between different locations, because the velocity field is arranged in “bulk flows” on $\sim 100$ Mpc scales. The form of these covariances depends on the type of velocity-field reconstruction applied (2024OJAp....7E..87B), and will be discussed further in Sec. 4. These correlations have potentially important implications for joint analyses of multiple standard sirens.
@@ 3 Determining the $H_{0}$ posterior
$\begin{split}&P(\mathrm{u}\,|\,H_{0},z,D,\mathrm{v})\propto\\ &\exp{\left[-\frac{\left(c\,z-H_{0}\,D-(r\,\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}})\,\mathrm{v}\right)^{2}}{(1-r^{2})\,\sigma_{u}^{2}}\right]}.\end{split}$ (5)
@@
and the covariance of the velocity differences $\mathbf{C}_{\mathrm{w}}$ is evaluated using Eq. 32 in 2024OJAp....7E..87B, and will be described in more detail in Sec. 4 for the velocity reconstruction method we adopt in our current study.
Eq. 6 can always be evaluated by numerical techniques which integrate over the multi-dimensional distance space, for any form of distance prior $P(D)$. However, if we represent the luminosity distance probability distributions $P(D_{\alpha})$ as Gaussian distributions with mean $\mu_{D,\alpha}$ and standard deviation $\sigma_{D,\alpha}$, Eq. 6 can be conveniently evaluated by analytical marginalisation techniques (2002MNRAS.335.1193B; 2010MNRAS.408..865T). Following the notation of 2010MNRAS.408..865T, we expand the log-likelihood $\mathcal{L}=-2\ln{P(\mathrm{u}_{\alpha})}$ of the inferred peculiar velocity field values as a Taylor series as a function of $\delta D_{\alpha}=D_{\alpha}-\mu_{D,\alpha}$,
@@
The marginalised log-likelihood $\mathcal{L}_{\rm marg}=-2\ln{L_{\rm marg}}$, after integrating over the Gaussian priors in $D_{\alpha}$, is then given by Eq. 15 in 2010MNRAS.408..865T,
@@ 4 Simulation
We tested our formalism for $H_{0}$ recovery by constructing a model bright siren analysis of a simulated dataset. For this purpose we used the $z=0$ snapshot of the Gigaparsec WiggleZ N-body (GiggleZ) simulation (2015MNRAS.449.1454P), which is a dark matter-only simulation run in a $1\,h^{-1}$ Gpc periodic box with a halo mass resolution of $3.0\times 10^{11}h^{-1}M_{\odot}$. This simulation was created using the fiducial WMAP-5 cosmological parameters: $\Omega_{m}=0.273$, $\Omega_{\Lambda}=0.727$, $\Omega_{b}=0.0456$, $h=0.705$, $\sigma_{8}=0.812$ and $n_{s}=0.960$.
@@
  • We considered two options for assigning distance errors to each model siren. In the first case, we added a Gaussian error to the distances equal to a constant fraction of the true distances, $\sigma_{D,\alpha}=f_{D}\,D_{{\rm true},\alpha}$, forming the noisy distances, $D_{\alpha}$. In the second option, we scaled the error with distance more realistically, by adopting the model of 2019PhRvD.100j3523M. Specifically, we scaled the signal-to-noise ratio (SNR) of the GW strain as $\rho_{\alpha}=\rho_{\rm min}D_{\rm max}/D_{{\rm true},\alpha}$ where $\rho_{\rm min}=12$, and the resulting error as $\sigma_{D,\alpha}=D_{{\rm true},\alpha}/\rho_{\alpha}$. We refer to this second case as “SNR-scaled distance errors”.
@@
We repeated this analysis varying the number of bright sirens $N$, their maximum distance $D_{\rm max}$, and the fractional distance error $f_{D}$ (or, as an alternative, the SNR-scaled errors).
For convenience of this analysis we assume a “Fourier-based” linear-theory velocity reconstruction approach, which uses a density-field defined within a Fourier cuboid. 2024OJAp....7E..87B contrast the Fourier-based approach with the more computationally-intensive spherical Fourier-Bessel reconstruction within a curved-sky window function, demonstrating in each case how the sample variance and noise propagates into the statistics of the reconstructed velocity field. However, in our study we focus instead on the propagation of the velocity field statistics into the posterior probability distribution of $H_{0}$. In this case, we simply assume a uniform selection function within the box.
@@
In this study we focus on the radial velocity of each source relative to an observer at the origin, $\mathrm{u}(\mathbf{x})=\sum_{i}\mathrm{u}_{i}(\mathbf{x})\,x_{i}/x$. The correlation between the radial velocities at two locations is given from Eq. 13 by (e.g., 1988ApJ...332L...7G; 2024MNRAS.527..501B),
@@
where $j_{\ell}$ are the spherical Bessel functions of order $\ell$ and $P_{m}(k)$ is the matter power spectrum, which we obtain for the GiggleZ fiducial cosmology using the CAMB and halofit prescriptions (2000ApJ...538..473L; 2003MNRAS.341.1311S; 2012ApJ...761..152T).
@@
Following 2024OJAp....7E..87B, the covariance of the reconstructed velocities between two positions, $\langle\mathrm{v}(\mathbf{x})\,\mathrm{v}(\mathbf{y})\rangle$, is obtained using the same relation given in Eq. 14 whilst making the replacement,
@@
These covariances allow us to determine the covariance of velocity differences $\mathbf{C}_{\mathrm{w}}$ introduced in Eq. 7, using Eq. 32 in 2024OJAp....7E..87B:
@@
where the factor $R=r\sigma_{\mathrm{u}}/\sigma_{\mathrm{v}}$ and $\epsilon_{\mathrm{u}}$ is the additional noise in the underlying velocities due to, for example, non-linear effects. We assume $\epsilon_{\mathrm{u}}=150\,{\rm km}\,{\rm s}^{-1}$ for the purposes of this study, which is a standard choice for residual velocity noise (2015MNRAS.450..317C; 2018ApJ...859..101S). Hence, we distinguish two contributions to the velocity errors: linear-theory covariance (due to the imperfection of velocity reconstruction, represented by the first four terms of Eq. 17) and non-linear noise (represented by the final term). For our fiducial assumptions the values of the coefficients are given by $\sigma_{u}=337\,{\rm km}\,{\rm s}^{-1}$, $\sigma_{v}=220\,{\rm km}\,{\rm s}^{-1}$ and $r=0.72$.
@@
In the following section we compare the results of applying these different analysis approaches for various choices of $N$, the number of sirens, $D_{\rm max}$, the maximum observed distance, and $f_{D}$, the fractional error in those distance measurements. We considered $N=[2,5,10,25,100]$, $D_{\rm max}=[50,100,150,200]$ Mpc, and $f_{D}=[0.01,0.02,0.05,0.10]$ to define this parameter space, adding the SNR-scaled errors as an alternative to $f_{D}$, making for 100 total tests. We repeated each of these tests for 50 different sub-samples drawn from the N-body simulation, saving the mean and standard deviation obtained from the $H_{0}$ posterior for each realisation. We then use the mean of these sets of results for a final measurement for each configuration.
Whilst the focus of our study is not to create realistic GW simulations, our studied configurations span a range of potential future GW datasets. The “detection power” of GW observatories is typically characterised by the distance range for detecting a Binary Neutron Star merger, equivalent to our $D_{\rm max}$ parameter. This range is currently $\approx 150$ Mpc for LIGO [^1] https://observing.docs.ligo.org/plan, although substantially lower for the Virgo ($\approx 50$ Mpc) and KAGRA ($\approx 10$ Mpc) GW observatories, which may be relevant because the identification of bright sirens with electromagnetic counterparts depends partly on localisation. The first bright siren event, GW170817, is located at a distance $\approx 44$ Mpc, with an initial distance error of $\approx 15\%$ (2017Natur.551...85A), which has been refined to $5-10\%$ by follow-up studies (e.g., 2019NatAs...3..940H; 2020MNRAS.492.3803H; 2022Natur.610..273M; KellyPaper).